OpenStudy (anonymous):

For anyone who's adventurous enough.... If a is a complex number and \[x ^{10}+x ^{5}+1=0\] Find \[x ^{2005}+1/x ^{2005}\]

OpenStudy (anonymous):

Also, second line should read "If x is a complex number"...sorry about the typo.

OpenStudy (jamesj):

hence \[ x^5 = \frac{-1 \pm \sqrt{3} i}{2} = e^{\pm \pi i/6} \]

OpenStudy (jamesj):

Now it becomes very easy to calculate x^2005 and x^-2005

OpenStudy (jamesj):

oops, wrong argument for exp: \[ x = e^{2\pi i /3} \ \ \ or \ \ \ x = e^{4 \pi i /3} \]

OpenStudy (anonymous):

So what are you getting as a final answer?....My approach is rather different from yours, but maybe both work?

OpenStudy (jamesj):

I'd rather not calculate it, but I know it's a real number.

OpenStudy (anonymous):

A real number, yes...and a fairly simple one at that...I'll post my solution in a few minutes...

OpenStudy (anonymous):

Here's my approach, which I think is much simpler... If we multiply the original equation by \[x ^{5}\] We get \[x ^{15}+x ^{10}+x ^{5}=0\] If we subtract the original equation from the above equation we get \[x ^{15}=1\] Since x cannot be 0 (as it would yield 1=0 in the original equation), dividing \[x ^{10}+x ^{5}+1=0 \] by \[x ^{5}\] yields \[x ^{5}+1/x ^{5}=-1\] and squaring that gets \[x ^{10}+1/x ^{10}=-1\] Now, \[x ^{2005}+1/x ^{2005}=(x ^{15})^{133}x ^{10}+1/(x ^{15})^{133}x ^{10}\] =\[x ^{10}+1/x ^{10}=-1\] Your way will likely give the right answer as well. Medal for your attempt.

OpenStudy (jamesj):

Not "likely", will. :-)